Grade 12Physical Chemistry

Find (a) the total number and (b) the total mass of neutrons in 7 mg of 14C. (Assume that mass of a neutron = 1.675 × 10–27 kg).

sudhanshu
12 Years agoGrade 12

1 Answer

Gaurav
11 Years ago
Number of atoms of 14C in 1 mole= 6.023 × 1023
Since 1 atom of 14C contains (14 – 6) i.e., 8 neutrons,
the number of neutrons in 14 g of 14C is (6.023 × 1023) ×8.
Or, 14 g of 14C contains (6.022 × 1023 × 8) neutrons.
Number of neutrons in 7 mg = 2.4092 × 1021
(b)
Mass of one neutron = 1.67493 × 10–27 kg
Mass of total neutrons in 7 g of 14C = (2.4092 × 1021) (1.67493 × 10–27 kg) = 4.0352 × 10–6 kg